1000cm³ of gas are converted at 27°c and 700mm of maccury. Convert this volume to s.t.p


1
Expert's answer
2021-02-28T07:23:27-0500

STP refer to 0oC (273 K) and a pressure of 760 mmHg.

27oC = 300 K

According to the combined gas law,


"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}"


Therefore,

"V_2=\frac{p_1V_1T_2}{T_1p_2}=\frac{700mmHg\times1000cm^3\times273K}{300K\times760mmHg}=838cm^3"


Answer: 838 cm3